NAOJ GW Elog Logbook 3.2

Today, the parts for upgrading the cryostat came. In particular, we have now an additional HV valve and an adapter for connecting KF25 endparts to a hose having 9mm diameter.
Still, a suitable hose is missing but I will order it Tomorrow.
I have successfully installed the items on the cryostat and are doing now test run for their vacuum performance.
Once the hose is also here, I will start the cryogenic test with He gas.

Here are now the final results on the simulations regarding the influence of the recoil mass of the PR, SR, and BS mirrors on the light scattering and the strain noise of KAGRA.
Final means that I will use these data for the publication which I am now writing.
I distinguished two different cases.
First, I assume a recoil mass made of a perfect lambertian scatterer. And second is to run the simulations with the BRDF data of titanium (roughly polished) which will be closest to reality.
The two last pictures are for BS, backside.
It should be noted that the data in the figures are given in 1/(sqrt(Hz) sr) and have to be multiplied by the respective solid-angle toward the beams waist to be comparable with KAGRAs goal sensitivity!
This will further decrease the values (by a magnitude of 10, or so), so I decided to leave the data as they are to keep a good overview of their relation to KAGRAs goal sensitivity.
It can be seen that even for the BS, we only have a neglible effect on the strain noise due to the scattering on the recoil mass, which is of course a good sign.

It would be nice if you check the polarization of the incident lights; maybe it would be different between the JASMINE one and the backscattering measurement system.

I measured the absorption at many modulation frequencies and plot the results in logaritmic scale.
I measured:
- Sapphire sample
- Reference sample: Bulk
- Reference sample: Surface

I present here the results of my own measurements on the polished SiC sample from Kyocera.
The measurements were done after Iwata-san did his measurements and after I did a necessary readjustment of the hight and the orientation of the laser.
The data show now a relatively equal maximum due to specular reflection, limited, of course, by the angular resolution of the system (-> 1 deg).
I added some roughly read data from Iwata-sans Backscattering measurements.
It is interesting to note that they are for all angles bigger than the JASMINE data would imply. At least for the maximum at AOI=0, it might be a true value. However, I am skeptical that the other values (all at around 0.02) are due to pure backscattering. Actually, I guess that noise is the main factor for their values...
The next step will be the measurement of the non-polished side of SiC.
It would be nice if you check the polarization of the incident lights; maybe it would be different between the JASMINE one and the backscattering measurement system.

I coded a Matlab script to evalue the analitic solution of the temperature distribution reported on the paper of Jackson et al.
I show the plot of T along z (depth) and r (radius)
Laser power = 1W
Surface:
absorption = 12ppm;
coating thickness = 10micron;
Bulk:
absorption = 20ppm/cm
thickness = 5mm;

Befor mesuring the Sapphire sample I made once again the calibration of the bulk reference sample (Schott glass#12 Abs=116%/cm) with low power 55mW and I found a calibrtion factor R=0.43W-1 (about 10% smaller than usually).
I measure the absorption of the Sapphire sample.
Pump power = 9.5W (maximum avaulable)
The scan along the Z axis Figure1 and Figure2 shows that the absorption value is 18 ppm/cm. The value has to be taken at the lowest inner point which is at Z=6mm in this case.
Then I moved the sample to that Z position and I made a map of 20x20mm with a resolution of 2x2mm.
I made it twice to check the repeatability (Figure3 and Figure4 ), and then I plot the difference Figure5 which is about the noise level: 1ppm/cm,
I made a higher resolution map 1x1mm (See Figure6).

An OSEM test bench in the ATC ISO-1 clean room is assembled by ATC people (Ikenoue-san and Saito-san).

These are now the results of the second round for the Ti sample. I rotated the sample by 90 degrees and measured its scattering again.
As can be seen, the peaks are much more sharper and more regular/symmetric compared to the non-rotated sample.
The strange shape of the AOI=0 curve is due to the interpolation process of the data gap (PD cuts off the Laser light) and the fact that I used one data point at theta=0 from older backscattering measurements.I need the interpolation for the implementation of the data into LightTools.

About BS oplev, its optical axis will be vertical, so it is relatively difficult for adults; maybe easier for japanese elementary school students.
One notice on this system is the angle of the axis is 37 degrees. Newport is selling adapters for 45 degrees for their mirror holders, which has a range of +/-7degrees, and that means even the combination of the adapter and the mirror holder cannot cover 37 degrees!!!!
Hmm, well, maybe I'll design a stuff...

Since bulk sample from LMA are rectangular, we needed a rectangular mount. We ordered the parts to be assembled and, after some trouble with metric and imperial lengths and screw threads, we succeded attach the magnets at the mount properly. (See Figure1 and Figure2 )
The sample of suprasil312 has a thickness of 20mm and a nominal absorption of 1.5-1.6ppm/cm.
I used the maximum power available setting the LD current at 7.5A. The measured power at the end of the pump path is 9.3W. I used the calibration factor R = 0.51 1/W which was measured on the high absorbing reference sample.
The scan along the z axis shows that the first surface is not absorbing but the second surface gives a big signal (600ppm). See Figure3 and Figure4.
Figure5 is a zoom of figure4 and it show the absorption inside the bulk (between 2 and 15mm) which has a value of 4ppm, while the noise outside the bulk is about 1ppm. This means that the measured absorption is 3ppm. About the double of the nominal value.

Members: Tatsumi, Hirose, Manuel
We went to Kashiwa campus on wednesday and we glued the wire breakers and the flags on the spare mirror.
I attach some pictures:
Figure1: Pour the first contact on the surface
Figure2: Spread the first contact
Figure3: Attach the little mesh to later remove the film.
Figure4: Set the mirror on the glueing rotating table. The arrow indicates the HR surface facing downward.
Figure5: Set the stages
Figure6: Set the lenses and rotate the table to be sure the white lines are aligned. Then remove the lenses.
Figure7: Put glue on the bigger wire breaker
Figure8: Put glue on the smaller wire breaker
Figure9: Fix the wire breakers on the stage
Figure10: Approach and attach the wire breakers to the surface using the micrometer screw
Figure11: Detail of the wire breaker
Figure12: Set the upper table and fix the flag mounts without glue to be sure they contact the surface
Figure13: Put glue on the flags
Figure14: Paste the flags
Figure15: All flags are set

Just for a small experiment: wash the loctite vac seal with acetone in a ultrasonic bath.
The attached pictures clearly show that the vac seal is broken to be rough.
Well, the kaptoned flexi circuit survives, and the micro D-sub connector also survives.
The soldering before this washing survives. The flux appears washed out by acetone; maybe the same effect can be expected with ethanol.
A short notice: I confirmed the LOCTITE vac seal can survive with ethanol in the ultrasonic bath.

I measured the scattering profiles of an unpolished Ti sample with JASMINEs scatterometer today.
The results in form of the calculated BRDF can be seen in the attached figure.
Also attached is a photograph of the Ti sample.

In order to state which part of the sample I was watching exactly, I used the sample boundaries as reference.
First I made a map with the DC signal in order to define the boundaries. I put it together in the same plot of the last absorption measurement (See Figure1.png). I could do that because I didn't remove the sample from the holder since the last measurement.
Then I took many pictures at the microscope and I put them together to make all mirror map. (See Figure2.jpg)
I overlap the two images making a matching of the boundaries. (See Figure3.jpg)
Finally I found the corresponding area. (See Figure4.jpg). The problem is that the pattern of dots at the microscope doesn't completely correspond to the absorption peaks pattern.

Members: Tatsumi, Fujii, Manuel.
JIG for gluing and necessary tools were delivered at NAOJ.
we wiped the tools one by one with ethanol before put inside the clean room. picture1.png
In order to carefully check the mechanical parts before gluing, we made a simulation of many of the gluing procedure steps:
- Unbox the JIG and clamp the basement on the optical table and attach the plastic move rod on the turn-table. picture2.png
- Set the alignment parts (the micrometer screws and the lenses) on the basement. picture3.png
- Through each lense we can see the magnified white mark to be aligned to the proper position of the mirror. picture4.png
- Remove the lenses and place the glueing parts for wire breakers: picture5.png and picture6.png
- Set the six posts to place the inner plate (3 shorter and 3 longer). And then, place the inner plate. picture7.png
(Unfortunately we found a little mistake on the hole position of one of the 3 shorter posts, we ask the company to fix it)
- Set the cilinder and close the JIG with the upper plate. picture8.png
- Open again the JIG and set the basement to stand up the JIG. picture9.png

We only have had 2D drawings of the mirror box's parts, and that prevents most of us from understanding the installation/hanging procedure of mirrors.
After the today's dry-run of the review (internal review) on the procedure, I've reconstructed the 3D data briefly from the 2D drawings, only which the manufacture company provided us (uploaded to the jgwdoc by Tatsumi-san).
If I didn't misunderstand the 2D drawings, the reconstructed 3D model shows a mechanical intereferance between one of the blue pillars and the mirror base (see the third figure attached).
And also, I think the base has a too narrow foot and should be widened!; the rail on the other side (the mirror-hanging jig) should be modified accrodingly.

In the last week I did some surface measurements with Zygos NewView8000 microscope on samples of Sapphire, SiO2-glass, and a GaAs-wafer from LMA.
Attached to this file are figures of the surface profile of each sample taken from the center and an area close to the edge of each sample.
The calculated surface roughnesses (rms) are as follows:
Sapphire-center: 1 nm
Sapphire-edge: 2nm
SiO2-center: 1nm
SiO2-edge: 1nm
GaAs-center: 2nm
GaAs-edge: 3nm

I made again a map of the same area of the LMA sample 15034. In the same conditions. I didn't unmount the sample from the holder, but I noticed that after one week it was dusty. So I cleaned it again with the first contact polymer. I show the picture of before and after cleaning. I show the comparison of the same measurement in the same conditions but after a week and a cleaning.

Fabian, Hirata, Shoda
We started the IM installation today.
Since we do not have a clamping tool for the cabling and some screws now, we are on the middle of the way.
I started to write an installation document:
https://docs.google.com/document/d/15Af4unTnrRd68tULCNv77kIi7h5c7p9246sUuCP1XCo/edit?usp=sharing
You can edit the document. Please read and write down if you find something missing.